Determine the molar solubility of PbSO4 in pure water. Ksp (PbSO4) = 1.6 × 10-10.

a) 4.12 × 10-5 M

b) 1.35 × 10-4 M

c) 4.89 × 10-4 M

d) 1.26 × 10-5 M

e) 1.28 × 10-8 M
Answer #1

Lets see the equilibrium reaction of PbSO4 to calculate molar solubility

PbSO4  \rightleftharpoons Pb2+(aq) + 2SO42-(aq)  

As we can see , every one mole of PbCl2 produces 1 mole of Pb2+ and 1 mole of SO4-

Lets see the ICE table

PbSO4\rightleftharpoons Pb2+(aq) + SO42-(aq)  

Initial - 0 0

Change - +s +s

at Equilibrium - s s

So, Solubility product constant for above reaction is

Ksp = [Pb2+][SO42-] = (s) (s) = s2

Ksp = s2

Then

s2 = Ksp = 1.6 \times 10-10

lets take square root on both the sides

we get

s = 1.26 \times 10-5 M

So option (d) is correct